OpenStudy (anonymous):

Which of the following is the simplified version of 1/2 log3 x – 2 log3 y + 4 log3 z ?

OpenStudy (campbell_st):

it can be written as \[\log _{3}(x^{1/2} y^{-1} z^4 ) \] or \[\log _{3}(\frac{x^{\frac{1}{2}}z^4}{y^2})\]